Market demand for a good is Q=240-3.5P. There are 2 firms on d market and their average cost are constant and 15. What quantity will each produce if they compete according to Betrands Model


1
Expert's answer
2021-10-18T09:07:28-0400

Average cost"=\frac{cost}{quantity}"

"Q=240-3.5P\\3.5P=240-Q\\P=68.57-0.286Q"

Average cost"=\frac{cost}{quantity}"

"15=\frac{68.57-0.286Q}{Q}\\15Q=68.57-0.286Q\\15Q+0.286Q=68.57\\15.286Q=68.57\\Q=4.49"